OpenStudy (anonymous):

What is the solution of the following system? {-2x-y=1 -4x-2y=-1 infinitely many solutions no solutions (3,8) (-3,-8) Will give a medal to whoever can help!

OpenStudy (anonymous):

@Isaiah.Feynman could you help?

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@JuanitaM

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

i'm not sure where to begin on this. usually on system problems there is a table or graph or something but there isn't on this one.

OpenStudy (anonymous):

make y the subject of the formula first, ill guide you as you do it :)

OpenStudy (anonymous):

y the subject of the formula for each equation first*

OpenStudy (anonymous):

would it be y=1-2x? is that what you mean?

OpenStudy (anonymous):

-2x-y=1 -2x = 1 + y y = 12x - 1 yes that's wat i ment :) now do the same for the other equation

OpenStudy (anonymous):

y = -2x - 1 sry my bad lol

OpenStudy (anonymous):

now do the same for -4x-2y=-1

OpenStudy (anonymous):

2y=-4x-1?

OpenStudy (anonymous):

u made an error, check over your signs

OpenStudy (anonymous):

-2y=-4x-1

OpenStudy (anonymous):

-4x-2y=-1 -4x = -1 + 2y 2y =-4x + 1 its positive 1 and not negative one

OpenStudy (anonymous):

when your making y the subject of the formula it must stand alone on its side of the equal sign like this y = -2x - 1. Therefore i want you to move the 2 next to the y in 2y =-4x + 1 to the other side of the formula

OpenStudy (anonymous):

of the equation, let me know wat you get :)